OpenStudy (tia_sweetheart17):

0.25-0.02(x+3)=-0.03(2-x)

OpenStudy (anonymous):

0.25-0.02x+(0.02)(3)=(-0.03)(2)-(-0.03)(x) 0.31-0.02x=-0.06+0.03x 0.31=0.04x-0.06 0.37=0.04x x=9.25